A boat takes 2/3 as much time to travel downstream as to return. If the rate of the river's current is 8 kph, what is the rate of the boat in still water?
Answer by TimothyLamb(4379)

time to return: t
time to go: (2/3)t
---
w = 8 kph
---
s = d / t
d = s * t
D = trip distance one way
---
D = (b + w)(2/3)t
D = (b - w)t
---
(b + w)(2/3)t = (b - w)t
(b + w)(2/3) = (b - w)
(2/3)b + (2/3)w = b - w
(2/3)w + (3/3)w = (3/3)b - (2/3)b
(5/3)w = (1/3)b
b = 3(5/3)8
---
b = 40 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php